OpenStudy (anonymous):

x=2y+3z------ solve for y ( i actually tried this one already and had an #epic fail)

jimthompson5910 (jim_thompson5910):

\[\large x=2y+3z\] \[\large x-3z=2y\] \[\large \frac{x-3z}{2}=y\] \[\large y=\frac{x-3z}{2}\] If you can, point out where you went wrong.

OpenStudy (anonymous):

On the second part!!! im soooo dumb!!!

jimthompson5910 (jim_thompson5910):

on the second step, I subtracted 3z from both sides because I want to isolate the y term. So I basically want y on one side while everything else is on the other.

jimthompson5910 (jim_thompson5910):

hopefully that clears things up a bit

OpenStudy (anonymous):

Jim im having trouble on # 44...

jimthompson5910 (jim_thompson5910):

sorry don't know what #44 is, is it posted somewhere on here?

OpenStudy (anonymous):

2ab+4=d----solve for a

OpenStudy (anonymous):

wana see wat i wrote?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

2ab+4=d -4 -4 2ab/4d ?????????????????? i suck major popsickles

jimthompson5910 (jim_thompson5910):

You're off to a great start \[\large 2ab+4=d\] \[\large 2ab=d-4\] \[\large a=\frac{d-4}{2b}\] In the last step I divided both sides by 2b (so the 2 and the b on the left side will cancel out and go away)

OpenStudy (anonymous):

Omg,,, i am on problem 60,,,, but im stuck!!! ugh its d(a-b)=c-------------- solve for a

jimthompson5910 (jim_thompson5910):

\[\large d(a-b)=c\] \[\large a-b=\frac{c}{d}\] \[\large a=\frac{c}{d}+b\]